A) What mass of oxygen must be used for complete combustion of 1 kg of an oil fuel with a 87% carbon and 13% hydrogen content?

B) What mass of flue gas would be produced for 1 kg of fuel completely combusted?
1
Expert's answer
2019-10-26T02:58:40-0400

First, the empirical formula of an oil fuel must be determined.

100 g of an oil fuel contains 87 g of carbon and 13 g of hydrogen. The number of moles of each element in an oil fuel equals:

n(C) = 87 g / 12 g/mol = 7.25 mol

n(H) = 13 g / 1g/mol = 13 mol

From here:

n(C) = 7.25 / 7.25 = 1

n(H) = 13 /7.25 = 1.8

The number must be multiplied by 5 to get integers:

n(C) = 5

n(H) = 9

As a result, the empirical formula of the compound is C5H9.

The combustion of the compound can be represented by the following equation:

4C5H9 + 9O2 = 20CO2 + 18H2O


A. Mass of oxygen must be used for complete combustion of an oil fuel with the empirical formula C5H9 equals:

m(O2) = (1000 g × 9 × 32 g/mol) / (4 × 69 g/mol) = 1043 g = 1.043 kg


B. Mass of CO2 produced from the combustion of an oil fuel with the empirical formula C5H9 equals:

m(CO2) = (1000 g × 20 × 44 g/mol) / (4 × 69 g/mol) = 3188 g = 3.188 kg


Answer: A. 1.043 kg of oxygen; B. 3.188 kg of CO2.